Find and correct the errors in the following mathematical statements.
4(x–5) = 4x–5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

4(x- 5)= 4x– 20 ≠4x–5 = RHS
The correct statement is 4(x-5) = 4x–20

Thanks